OpenStudy (anonymous):

Let f(x) = 49 - x2 , g(x) = 7 Find (f-g)(X) A. -x^2 - x + 56 B. 7 + x C. -x^2 + x + 42 D. -x^3 - 7x^2 - 49x + 343

OpenStudy (anonymous):

(f - g)(x) = f(x) - g(x) = 49 - x^2 - 7 = -x^2 + 42

OpenStudy (anonymous):

Its g(x) = 7 - x forgot to put the - x

OpenStudy (anonymous):

f(x) - g(x) = 49 - x^2 - (7 - x) = 49 - x^2 - 7 + x = -x^2 + x + 42

OpenStudy (anonymous):

(f - g)(x) = f(x) - g(x) Likewise, (f + g)(x) = f(x) + g(x)

OpenStudy (anonymous):

Yes, C is the answer

OpenStudy (anonymous):

sweet thanks alot preciate it